OpenStudy (anonymous):

f(x)=2x+4; find f(3x)-3f(x)

OpenStudy (jdoe0001):

\(\bf f(x)=2x+4\\\\ \quad \\ f(3x) = \color{red}{2(3x)+4}\\ 3f(x) = \color{blue}{3 \times (2x+4)}\\ \quad \\ \quad \\ f(3x)-3f(x) = [\color{red}{2(3x)+4}] \quad - \quad [\color{blue}{3 \times (2x+4)}]\)

OpenStudy (anonymous):

i tried pluging them in but i'm gettting confused

OpenStudy (jdoe0001):

\(\bf f(x)=2x+4\\\\ \quad \\ f(3x) = 2(3x)+4\implies 6x+4\\ 3f(x) = 3 \times (2x+4)\implies 6x+12\\ \quad \\ \quad \\ f(3x)-3f(x) = [6x+4] \quad - \quad [6x+12]\)

OpenStudy (jdoe0001):

-[ 6x + 12] = -6x - 12x

OpenStudy (jdoe0001):

what would that give you?

OpenStudy (anonymous):

-6x-12

OpenStudy (jdoe0001):

\(\bf f(3x)-3f(x) = [6x+4] \quad - \quad [6x+12]\implies \cancel{6x}+4\cancel{-6x}-12\)

OpenStudy (anonymous):

-8

OpenStudy (anonymous):

see where i messed up was i got 6x-fx+4 don't know how i did that

OpenStudy (jdoe0001):

it happens, tis ok :)